A company typically receives an order for a particular product once every five days.

a. What is the probability that there are two orders in 12 days?

b. In 30 days?

c. In one day?
Answer #1

Answer)

Answer)

Here we need to use the binomial formula

P(r) = ncr*(p^r)*(1-p)^n-r

Ncr = n!/(r!*(n-r)!)

N! = N*n-1*n-2*n-3*n-4*n-5........till 1

For example 5! = 5*4*3*2*1

Special case is 0! = 1

A)

P = probability of single trial = 1/5 = 0.2

N = number of trials = 12

R = desired success = 2

After substitution

Required probability is = 0.28346784154

B)

N = 30

Rest is same

P(2) = 0.03365649482

C)

Here probability is 0, as probability of getting an order is once in every 5 days

So getting two orders in one day is highly unlikely and probability will be close to 0